# Symbolic Mixture-of-Experts: Adaptive Skill-based Routing for Heterogeneous Reasoning

### v2

@@ 3.2.1 Model Profile Creation
To recruit the $k$ most suitable experts for a given question, we assess each model’s specialized skills across various problem-solving domains, illustrated in Fig. 2 (a). This is done by evaluating their performance on the validation set for each task (see Table 6 for sizes), thereby constructing a model profile $P_{i}$ for each model $M_{i}$. For each question in the validation set, we first prompt an LLM -- referred to as the ‘‘Keyword LLM’’ -- to identify the essential skills required to solve the problem. The prompt used for generating these required skills is provided in Appendix G. For consistency, we use Qwen2.5-7B-Instruct (Team, 2024) as the Keyword LLM. To reduce noise, we generate keyword annotations for each question five times, and retain only those that appear more than once across the whole validation set. These extracted skills represent core knowledge areas necessary for solving the problem -- for instance, a given college-level math problem may require skills such as algebra, calculus, and geometry. Once all questions are annotated with their required skills, each model $M_{i}$ in the pool attempts to solve them using Chain-of-Thought reasoning (Wei et al., 2022). A correct answer increases the score of each associated skill by $+1$, while an incorrect answer results in a $-1$ penalty. At the end of this process, each model has a profile $P_{i}\in\{P_{1},\dots,P_{n}\}$ represented as a dictionary. For example, a model’s profile may be: {‘Algebra’: 10, ‘Biology’: 3, ‘Chemistry’: -6, ...}.
@@ 3.2.2 Aggregator Selection
An aggregator is a model that consolidates $k$ outputs into a single high-quality response. Our pilot experiments, along with findings from Wang et al. (2024a) and Li et al. (2024b), indicate that the aggregator model plays a crucial role in the final performance, and selecting the most effective model for aggregation is a non-trivial challenge. We find that reasoning capability and aggregation ability are often orthogonal. That is, a strong reasoning model is not necessarily a strong aggregator and vice versa; qualitatively, we show this later in Table 5. We find that adaptively selecting an aggregator on the instance level based on model profiles is less effective, motivating us to choose the aggregator based on a model’s ability to consolidate different answers. To identify the best aggregator per task, we construct a synthetic task using the same validation set. From the profile creation process, we obtain outputs from all models, some correct and some incorrect. For each question, we sample one correct reasoning chain and two incorrect ones, structuring the input as follows: {question}, {correct_CoT}, {incorrect_CoT}, {incorrect_CoT}. We shuffle the order of the correct and incorrect CoTs and instruct each model to act as an aggregator (using the prompt shown in Appendix G), synthesizing a final output with a predicted answer. We then benchmark each model’s aggregation ability, measuring how well it can generate a correct answer based on this input, and select the best-performing aggregator for each task.
@@ 3.3.1 Skill-based Recruiting
At inference time (see Fig. 2 (b)), we follow the same keyword annotation procedure as in Section 3.2.1 to generate relevant keywords for the test sample. To align a test sample’s keywords with those in the model profiles, we employ Sentence-BERT (Reimers & Gurevych, 2020) to match keywords via the cosine similarity between their embeddings. Each test-time keyword is then mapped to its closest counterpart in the model profile. Next, the expert recruitment is performed by selecting the top $k$ models whose profiles best match the required skills of the test sample. This is determined by two factors: (1) local suitability score and (2) global competency. For each model $M_{i}$, its local suitability score for a test sample $q$, $\mathcal{S}(M_{i},q)$ is computed as the sum of its skill scores over the set of keywords needed for $q$ (denoted as $K_{q}$). It can be expressed as follows:
@@
where $s^{(i)}_{k_{j}}$ represents the score of model $M_{i}$ for the $j$-th skill in the test sample $q$. This results in an model ranking distribution $\mathcal{D}_{q}$ for each test sample $q$:
$\mathcal{D}_{q}=\{\mathcal{S}(M_{1},q),\mathcal{S}(M_{2},q),...,\mathcal{S}(M_{n},q)\}$
Intuitively, suppose $M_{1}$ has scores of $+3$, $+5$, and $-2$ for algebra, calculus, and geometry, respectively, which are needed for a given sample; its total score for this sample would be $3+5-2=6$. Having this calculation for all models yields a ranking distribution of model strengths for the given sample, e.g., {$M_{1}$: 6, $M_{2}$: 3, ..., $M_{n}$: -10}, which is the ranking of how suitable a model is for a sample. To account for each model’s overall strength in a task, i.e., global competency, we compute each model’s total score across all keywords in their profile, and normalize it by the total sum of all agents’ scores. We denote this global strength as $\gamma_{i}$, representing a model’s overall task performance relative to others. Finally, the expert selection is performed by sampling from the product of the local suitability score, $\mathcal{S}(M_{i},q)$ (from a model rank distribution $\mathcal{D}_{q}$) and the global competency $\gamma_{i}$. That is, the relevance score of a model $M_{i}$ for a test sample $q$ is:
$w^{(i)}_{q}=\gamma_{i}\mathcal{S}(M_{i},q)$
We apply a softmax function with the temperature set to $0.5$ to this distribution $\{w^{(i)}_{q}\}_{i=1}^{n}$, and then sample $k$ experts for each problem, i.e.,
@@
To enhance efficiency and reduce computational overhead, we filter out low-frequency experts -- those who appear in fewer than $5\%$ of test cases. For example, given a test set with 100 samples, where 3 experts are recruited per sample (totaling 300 selections), any expert appearing fewer than $300\times 5\%=6$ times is discarded and resampled from the remaining higher-frequency experts.
@@ 3.3.2 Final Answer Generation
After expert recruitment, each sample will be passed to the experts, i.e., the input for each expert is the test problem, $x_{0}=q$. These experts generate their reasoning paths to the problem in the form of Chain-of-Thought (Wei et al., 2022):
$y_{0}^{(i)}=E^{(i)}(x_{0})~{}\forall~{}i\in\{1,2,...,k\}$
@@ 4 Results and Analysis
4.1 Experimental Setup
Model Pool and Datasets.
Our experiments consider $16$ LLMs ranging from 3.5B to 12B parameters, with most models falling in the 7–8B range. These include general-purpose instruction-tuned models, domain-specific fine-tuned variants on math and biology, and models distilled from DeepSeek R1’s trajectories (DeepSeek-AI et al., 2025a). A full list of models is provided in Table 7 in the Appendix. We measure performance on a diverse range of datasets, chosen to require expertise in a number of domains. First, we consider MMLU-Pro (Wang et al., 2024c), which is a harder version of MMLU (Hendrycks et al., 2021), containing a variety of questions across 14 college-level subjects, from Philosophy and History to Economics and Chemistry. Given its large test set of 12,000 samples and the computational cost of evaluating proprietary models, we employ stratified sampling to create a subset of 2,100 samples, ensuring each category contains 150 samples. We also evaluate on AIME 2024, which is a challenging mathematics competition dataset containing math Olympiad problems. For more science-specific reasoning, we test on GPQA (Rein et al., 2023), which contains questions across a variety of science fields written by experts, explicitly designed to be difficult to answer even by skilled humans with web access. Finally, we also include MedMCQA (Pal et al., 2022), which covers questions sourced from medical entrance exams across 21 medical subjects. For each dataset, we sample around 350 samples as the validation set to create the agent and aggregator selection profiles. [^2] For AIME, we sample validation questions from prior years’ problems (2012-2023). Full dataset statistics are provided in Table 6 in the Appendix.
Baselines.
We compare against four categories of baselines.
  • Zero-shot single-model methods: This category includes proprietary models such as GPT-4o-mini (OpenAI, 2024), Gemini 1.5 Pro (Team et al., 2024), and DeepSeek-V3 (DeepSeek-AI et al., 2025b); high-capacity open-source models like Qwen2.5 72B (Qwen et al., 2025) and Llama 3.3 70B (AI@Meta, 2024); and strong distilled 7B models such as QwenR1 (DeepSeek-AI et al., 2025a). For reference, we also report the best task-specific model from our pool for each task, denoted as Task-Best.
For single-model baselines, we use the strongest task-specific LLM, while for multi-model baselines, we select the top three models per task. These selections, like those in Symbolic-MoE, are determined based on validation performance. Table 8 in the Appendix details the top-1 and top-3 models for each task.
Implementation Details.
We conduct our experiments for Symbolic-MoE and other single-model baselines on a single A6000 GPU with 48 GB of memory, while MoA and ReConcile are executed on 8 A6000 GPUs for parallelization. For the 70B models, we use the original version without quantization and perform inference on 4 A6000 GPUs. All open-source models utilize vLLM (Kwon et al., 2023) for inference. The temperature is set to 0.7 for all methods. The maximum output token length is fixed at $4096$ for all models, except for QwenR1 and LlamaR1, which have a limit of $32768$ since they are trained with longer trajectories and tend to generate longer outputs. All results, except those from proprietary models (due to budget constraints), are averaged over three random seeds. Further details on the model pool, distribution of the expert recruited, and all the prompts we use can be found in Table 7 and Appendix G.
@@ 4.2 Main Results
Across all domains, Symbolic-MoE shows superior performance compared to all baselines, beating single-model baselines using the best overall model (e.g., SR, SC), multi-agent debate with a single model (Debate and Self-MoA), as well as multi-model multi-agent baselines like MoA and ReConcile. Symbolic-MoE outperforms the most competitive multi-agent baseline, Self-MoA, by $8.15\%$ (absolute) on average, with consistent improvements across domains (e.g., $8.28\%$ gain on MMLU-Pro, $13.45\%$ on AIME, $4.92\%$ on GPQA, $6.08\%$ on MedMCQA). These gains are also seen when comparing to multi-model baselines like MoA and ReConcile that use the top three strongest models per domain. Symbolic-MoE also substantially outperforms test-time scaling methods, such as Self-Refine (Madaan et al., 2023b) and Self-Consistency (Wang et al., 2023a). Surprisingly, when using the task-best model, SC beats multi-agent debate baselines (e.g., Self-MoA, MoA), though it still underperforms Symbolic-MoE by an average of $3.71\%$. This indicates that scaling test-time compute with the task-best model is a simple but effective way to improve performance, and adaptively selecting the most suitable experts leads to further improvements.
Symbolic-MoE generalize well across tasks.
@@
No single baseline in Table 1 is universally effective across all tasks. For instance, while MoA performs well on MMLU-Pro, it struggles on AIME; ReConcile excels in MedMCQA but fails to generalize to GPQA. Knowing which method works best for a task is therefore nontrivial, requiring running every method on validation sets to choose from the many settings available. In contrast, Symbolic-MoE consistently delivers a strong performance across all domains. Symbolic-MoE especially excels on AIME and GPQA, where SC is a surprisingly strong baseline, and where other strong methods like MoA and Self-MoA fall far behind Symbolic-MoE. Moreover, we see that, while SC with the top model is the most competitive setting on AIME and GPQA, it lags behind other baselines on MMLU-Pro and MedMCQA, where multi-agent baselines perform better. This discrepancy may stem from the broader subject diversity in MMLU-Pro and MedMCQA, where agent collaboration facilitates better consensus, whereas AIME is more math-focused, and favors individual model performance -- the task-best model, QwenR1 7B, delivers strong solo performance already. In light of that, we include QwenR1 7B in Table 1, which is a powerful model distilled from DeepSeek-R1’s trajectories (DeepSeek-AI et al., 2025a). While QwenR1 demonstrates exceptional math and code reasoning capabilities ($55.93\%$ on AIME), leading to strong Self-Consistency performance on AIME ($67.78\%$), it struggles to generalize to other domains such as MedMCQA. This further underscores the need for a robust and flexible framework like Symbolic-MoE to achieve broad generalization across diverse tasks.
Symbolic-MoE matches strong proprietary models and larger 70B models.
@@
In Table 1, we also find that Symbolic-MoE achieves a similar average performance to models that have substantially more parameters. For example, Symbolic-MoE outperforms Llama3.3 70B on AIME and GPQA and roughly matches it on MedMCQA, despite requiring only four 7-8B models (three for the experts and one for the aggregator). Similarly, Symbolic-MoE outperforms or matches a number of strong proprietary models on average -- for instance, it matches Gemini 1.5 Pro and outperforms GPT4o-mini, driven in part by sizable gains on AIME and GPQA. These results underscore that, by drawing from a large pool of experts, Symbolic-MoE enables a smaller set of models to achieve performance comparable to models with significantly more parameters, especially when considering a heterogeneous set of tasks like the ones we explore.
@@
Table 2: Efficiency comparison of MoA and Symbolic-MoE. We report time per sample, based on the total time needed to perform inference for the whole GPQA test set (divided by the number of questions in the test set).
Symbolic-MoE is more efficient.
We compare Symbolic-MoE’s efficiency to a naive implementation of sequential inference and to that of MoA and in Table 2. On GPQA, we measure the average run-time for the 198 examples in the test set. Unsurprisingly, the sequential inference baseline shows the highest latency, since the model is constantly changing, requiring loading and offloading it for each instance. We also find that Symbolic-MoE operates in $44\%$ less time on a single GPU than MoA, which requires discussion, while achieving better accuracy. Also, Symbolic-MoE running on a single GPU shows similar time with MoA on 4 GPUs, and furthermore, when we run Symbolic-MoE on 4 GPUs, it results in almost 2×2$\times$2 × speedup over MoA.
@@
Table 3: Comparison of Symbolic-MoE with and without discussion across varying aggregators. Discussion stabilizes performance with suboptimal aggregators (first four rows) but has little effect with an optimal one (last two rows).
Symbolic-MoE achieves efficiency in two key ways: (1) it employs batched inference to minimize model loading time, and (2) it eliminates round-wise discussions, which introduce significant overhead at each round. In Table 3 we find that with a reliable aggregator, skipping inter-agent discussions yields performance comparable to actually conducting them. Engaging $k$ agents in a one-round discussion requires generating $k$ initial outputs, concatenating them, and then generating $k$ updated outputs, significantly increasing overhead.
Like multi-agent discussion baselines, Symbolic-MoE can also operate in a discussion-based manner. Instead of aggregating initial responses, models first engage in a round discussion before submitting final answers to an aggregator. Table 3 evaluates this approach, comparing the adaptive aggregator (suboptimal) and task-best aggregator (suboptimal), as well as the task-specific aggregator (optimal) on MMLU-Pro and GPQA. Given the optimal aggregator, we see that discussion yields marginal gains on MMLU-Pro ($63.83$ vs. $63.71$) and a modest drop on GPQA ($57.72$ vs. $57.78$). We find that while round-wise discussion does improve performance incrementally, but the final outcome is ultimately determined by the strength of the aggregator. As a result, Symbolic-MoE has less run time than the baselines while maintaining or even surpassing their performance, as demonstrated in Table 2.
4.3 Analysis
Utility of the Model Profile.

### v3

@@ 3.2.1 Model Profile Creation
To recruit the $k$ most suitable experts for a given question, we assess each model’s specialized skills across various problem-solving domains, illustrated in Fig. 2 (a). This is done by evaluating their performance on the validation set for each task (see Table 13 for sizes), thereby constructing a model profile $P_{i}$ for each model $M_{i}$. For each question in the validation set, we first prompt an LLM – referred to as the “Keyword LLM” – to identify the essential skills required to solve the problem. For consistency, we use Qwen2.5-7B-Instruct (Qwen Team, 2024) as the Keyword LLM. To reduce noise, we generate keyword annotations for each question five times, and retain only those that appear more than once across the whole validation set. These extracted skills represent core knowledge areas necessary for solving the problem – for instance, a given college-level math problem may require skills such as algebra, calculus, and geometry. Once all questions are annotated with their required skills, each model $M_{i}$ in the pool attempts to solve them using Chain-of-Thought reasoning (Wei et al., 2022). A correct answer increases the score of each associated skill by $+1$, while an incorrect answer results in a $-1$ penalty. At the end of this process, each model has a profile $P_{i}$ represented as a dictionary, e.g., {‘Algebra’: 10, ‘Biology’: 3, ‘Chemistry’: -6, …}.
@@ 3.2.2 Aggregator Selection
An aggregator is a model that consolidates $k$ outputs into a single high-quality response. Our pilot experiments, along with findings from Wang et al. (2024a) and Li et al. (2025), indicate that the aggregator model plays a crucial role in the final performance, and selecting the most effective model for aggregation is non-trivial. We find that a strong reasoning model is not necessarily a strong aggregator and vice versa; qualitatively, we show this later in Table 6. We also find that adaptively selecting an aggregator on the instance level based on model profiles is less effective, motivating us to choose the aggregator based on its ability to select correct answers. To identify the best aggregator per task, we construct a synthetic task using the same validation set. From the profile creation process, we obtain outputs from all models, some correct and some incorrect. For each question, we sample one correct reasoning chain and two incorrect ones, structuring the input as follows: {question}, {correct_CoT}, {incorrect_CoT}, {incorrect_CoT}. We shuffle the order of the CoTs and instruct each model to act as an aggregator (using the prompt shown in Appendix L), synthesizing a final output with a predicted answer. We then benchmark each model’s aggregation ability and select the best-performing aggregator for each task.
@@ 3.3.1 Skill-based Recruiting
At inference time (see Fig. 2 (b)), we follow the same keyword annotation procedure as in Section 3.2.1 to generate relevant keywords for the test sample. To align a test sample’s keywords with those in the model profiles, we employ Sentence-BERT (Reimers & Gurevych, 2020) to match keywords via the cosine similarity between their embeddings. Next, expert recruitment is performed by selecting the top $k$ models whose profiles best match the required skills of the test sample. This is determined by two factors: (1) local suitability score and (2) global competency. For each model $M_{i}$, its local suitability score for a test sample $q$, $\mathcal{S}(M_{i},q)$ is computed as the sum of its skill scores over the set of keywords needed for $q$ (denoted as $K_{q}$). It can be expressed as follows:
@@
where $s^{(i)}_{k_{j}}$ represents the score of model $M_{i}$ for the $j$-th skill in the test sample $q$. This results in an model ranking distribution $\mathcal{D}_{q}$ for each test sample $q$: $\mathcal{D}_{q}=\{\mathcal{S}(M_{1},q),\mathcal{S}(M_{2},q),...,\mathcal{S}(M_{n},q)\}$.
Intuitively, suppose $M_{1}$ has scores of $+3$, $+5$, and $-2$ for algebra, calculus, and geometry, respectively, which are needed for a given sample; its total score for this sample would be $3+5-2=6$. Calculating this score for all models yields a distribution of model strengths for the given sample, e.g., {$M_{1}$: 6, $M_{2}$: 3, …, $M_{n}$: -10}, which is the ranking of how suitable a model is for a sample.
To account for each model’s overall strength in a task, i.e., global competency, we compute each model’s total score across all keywords in their profile, and normalize it by the total sum of all agents’ scores. We denote this global strength as $\gamma_{i}$, representing a model’s overall task performance relative to others. Finally, the expert selection is performed by sampling from the product of the local suitability score, $\mathcal{S}(M_{i},q)$ (from a model rank distribution $\mathcal{D}_{q}$) and the global competency $\gamma_{i}$. That is, the relevance score of a model $M_{i}$ for a test sample $q$ is: $w^{(i)}_{q}=\gamma_{i}\mathcal{S}(M_{i},q)$. We apply a softmax function with the temperature set to $0.5$ to this distribution $\{w^{(i)}_{q}\}_{i=1}^{n}$, and then sample $k$ experts for each problem, i.e.,
@@
To enhance efficiency, we trim low-frequency experts, i.e., those who appear in fewer than $5\%$ of test cases. For example, given a test set with 100 samples, where 3 experts are recruited per sample (totaling 300 selections), any expert appearing fewer than $300\times 5\%=6$ times is discarded and resampled from the remaining higher-frequency experts. To visualize this, we provide the expert distribution before and after trimming in Fig. 4.
@@ 4 Results and Analysis
4.1 Experimental Setup
We consider $16$ LLMs ranging from 3.5B to 12B parameters, with most models falling in the 7–8B range. These include general-purpose instruction-tuned models, domain-specific fine-tuned variants on math and biology, and models distilled from DeepSeek R1’s trajectories (DeepSeek-AI et al., 2025a). A full list of models is provided in Table 9. We measure performance on a diverse range of datasets, chosen to require expertise in a number of domains. First, we consider MMLU-Pro (Wang et al., 2024c), which is a harder version of MMLU (Hendrycks et al., 2021a), containing a variety of questions across 14 college-level subjects. Given its large test set of 12,000 samples and the computational cost of evaluating proprietary models, we employ stratified sampling to create a subset of 2,100 samples, ensuring each category contains 150 samples. We also evaluate on AIME 2024, which is a challenging mathematics competition dataset containing math Olympiad problems. For more science-specific reasoning, we test on GPQA Diamond (Rein et al., 2023), which contains questions across a variety of science fields written by experts, explicitly designed to be difficult to answer even by skilled humans with web access. Finally, we include MedMCQA (Pal et al., 2022), which covers questions sourced from medical entrance exams across 21 medical subjects. For each dataset, we sample around 350 samples as the validation set to create the model profiles. [^2] For AIME, we sample validation questions from prior years’ problems (2012-2023). Full dataset statistics are provided in Table 13, and the model pool we consider is shown in Appendix D. For single-model baselines, we use the strongest task-specific LLM, while for multi-model baselines, we select the top three models per task. These selections, like those in Symbolic-MoE, are made based on validation performance, which is summarized in Table 10. More details on baselines and their implementations can be found in Appendix B and Appendix C.
@@ 4.2 Main Results
Symbolic-MoE consistently outperforms all baselines. Across all domains, Symbolic-MoE shows superior performance compared to all baselines, beating single-model baselines (e.g., SR, SC) using the best overall model, multi-agent debate with a single model (e.g., Debate, Self-MoA), as well as multi-model multi-agent baselines (e.g., MoA, ReConcile). Symbolic-MoE outperforms the most competitive multi-agent baseline, Self-MoA, by $8.15\%$ (absolute) on average, with consistent improvements across domains (e.g., $8.28\%$ gain on MMLU-Pro, $13.45\%$ on AIME, $4.92\%$ on GPQA, $6.08\%$ on MedMCQA). These gains are also seen when comparing to multi-model baselines like MoA and ReConcile that use the top three strongest models per domain. Symbolic-MoE also substantially outperforms test-time scaling methods, such as SR (Madaan et al., 2023b) and SC (Wang et al., 2023a). Surprisingly, with the task-best model, SC beats multi-agent debate baselines (e.g., Self-MoA, MoA), though it still underperforms Symbolic-MoE by an average of $3.71\%$. This indicates that scaling test-time compute with the task-best model is a simple yet effective way to improve performance, and adaptively selecting suitable experts leads to further improvements.
@@
Symbolic-MoE generalizes well across tasks. No single baseline in Table 1 is universally effective across all tasks. For instance, while MoA performs well on MMLU-Pro, it struggles on AIME; ReConcile excels in MedMCQA but fails to generalize to GPQA. Knowing which method works best for a task is therefore nontrivial, requiring running every method on validation sets to choose from the many settings available. In contrast, Symbolic-MoE consistently delivers strong performance across all domains. Symbolic-MoE especially excels on AIME and GPQA, where SC is a surprisingly strong baseline, and where other strong methods like MoA and Self-MoA fall far behind Symbolic-MoE. Moreover, we see that, while SC with the top model is the most competitive setting on AIME and GPQA, it lags behind other baselines on MMLU-Pro and MedMCQA, where multi-agent baselines perform better. This discrepancy may stem from the broader subject diversity in MMLU-Pro and MedMCQA, where agent collaboration facilitates better consensus, whereas AIME is more math-focused, and favors individual model performance – the task-best model, QwenR1 7B, delivers strong solo performance already. In light of that, we include QwenR1 7B in Table 1, which is a powerful model distilled from DeepSeek-R1’s trajectories (DeepSeek-AI et al., 2025a). While QwenR1 demonstrates exceptional math and code reasoning capabilities ($55.93\%$ on AIME), leading to strong Self-Consistency performance ($67.78\%$), it struggles to generalize to other domains such as MedMCQA. This further underscores the need for a robust and flexible framework like Symbolic-MoE to achieve broad generalization across diverse tasks.
@@
Symbolic-MoE matches strong proprietary models and larger 70B models. In Table 1, we also find that Symbolic-MoE achieves a similar average performance to models that have substantially more parameters. For example, Symbolic-MoE outperforms Llama3.3 70B on AIME and GPQA and roughly matches it on MedMCQA, despite requiring only four 7-8B models (three for the experts and one for the aggregator). Similarly, Symbolic-MoE outperforms or matches a number of strong proprietary models on average – for instance, it matches Gemini 1.5 Pro and outperforms GPT4o-mini, driven in part by sizable gains on AIME and GPQA. These results underscore that, by drawing from a large pool of experts, Symbolic-MoE enables a smaller set of models to achieve performance comparable to models with significantly more parameters, especially when considering a heterogeneous set of tasks like ours.
4.3 Efficiency Anaylsis
@@
Table 2: Efficiency comparison of MoA and Symbolic-MoE on the GPQA test set. Run time is averaged per sample.
Run time efficiency of Symbolic-MoE. We compare Symbolic-MoE’s efficiency to a naive implementation of sequential inference and to that of MoA and in Table 2. We measure the average run-time at inference on GPQA. Unsurprisingly, the sequential inference baseline shows the highest latency, since the model is constantly changing, requiring loading and offloading it for each instance. We also find that Symbolic-MoE operates in $44\%$ less time on a single GPU than MoA, which requires multi-round discussion, with better accuracy. Moreover, Symbolic-MoE running on a single GPU shows a similar time with MoA on 4 GPUs, and furthermore, when we run Symbolic-MoE on 4 GPUs, it results in almost 2×2$\times$2 × speedup over MoA. We provide another efficiency measurement, token count comparison, in Appendix H, where we find that Symbolic-MoE has the best trade-off between token usage and performance among multi-agent baselines.
@@
Table 3: Comparison of with and without discussion across varying aggregators. Discussion stabilizes performance with suboptimal aggregators but has little effect with an optimal aggregator.
Methodological efficiency of Symbolic-MoE. Like multi-agent discussion baselines, Symbolic-MoE can also operate in a discussion-based manner. Instead of aggregating initial responses, models first engage in a round discussion before submitting final answers to an aggregator. Table 3 evaluates this approach, comparing the adaptive aggregator (suboptimal) and task-best aggregator (suboptimal), as well as the task-specific aggregator (optimal) on MMLU-Pro and GPQA. Given the optimal aggregator, we see that discussion yields marginal gains on MMLU-Pro ($63.83$ vs. $63.71$) and a modest drop on GPQA ($57.72$ vs. $57.78$). We find that while round-wise discussion does improve performance incrementally, but the final outcome is ultimately determined by the strength of the aggregator. As a result, Symbolic-MoE has less run time than the baselines while maintaining or even surpassing their performance, as demonstrated in Table 2.
4.4 Additional Analysis